# Landau Theory of Barocaloric Plastic Crystals

### v3

@@ I Introduction
The discovery of colossal BC effects in PCs has sparked theoretical and computational interest primarily using density functional theory [24], molecular dynamics methods [25, 15, 26, 27], supercell lattice dynamics calculations [11], and mean-field microscopic modelling [28], aimed at understanding plastic-to-crystal transitions and predicting their associated BC response. However, a macroscopic description based on Landau theory is missing. The purpose of this work is to propose such a theory, and illustrate it with a mimimal model that captures the essential features of plastic-to-crystal phase transitions. The main advantage of this approach is that both the molecular and lattice symmetries are incorporated into the model by construction, as well as the couplings between the order parameter (OP) of the transition and the elastic strains. We consider the simplest non-trivial case in which the orientational ordering is described by a quadrupolar OP, as is often done for nematic ordering in liquid crystals. Consequently, our model draws parallels to the Landau-de Gennes theory of phase transitions in liquid crystals [29].
@@ II.2 Free-energy density
${\bm{\epsilon}}=-\frac{1}{3K}\left\{\left[-K\alpha\left(T-T_{0}\right)+P+\left(g-e/3\right)\operatorname{Tr}{\bm{Q}}^{2}\right]\mathbb{1}+d\mathbf{Q}+e{\bm{Q}}^{2}\right\}.$ (2)
@@
$\tilde{b}$ $=b-\frac{2de}{3K}\left(1-\frac{\mu}{3K}\right),$
$\tilde{c}$ $=c-\frac{1}{2K}\left[g^{2}+\frac{e^{2}}{9}\left(1-\frac{\mu}{3K}\right)\right],$
$\tilde{T}_{Q}$ $=\left[T_{Q}+\frac{d^{2}}{3Ka}\left(1-\frac{\mu}{3K}\right)+\frac{g\alpha}{a}T_{0}\right]\left[1+\frac{g\alpha}{a}\right]^{-1}.$

### v4

@@ I Introduction
The discovery of colossal BC effects in PCs has sparked theoretical and computational interest primarily using density functional theory [24], molecular dynamics methods [25, 15, 26, 27], supercell lattice dynamics calculations [11], and mean-field microscopic modelling [28], aimed at understanding plastic-to-crystal transitions and predicting their associated BC response. However, a macroscopic description based on Landau theory is missing. The purpose of this work is to propose such a theory, and illustrate it with a minimal model that captures the essential features of plastic-to-crystal phase transitions. The main advantage of this approach is that both the molecular and lattice symmetries are incorporated into the model by construction, as well as the couplings between the order parameter (OP) of the transition and the elastic strains. We consider the simplest non-trivial case in which the orientational ordering is described by a quadrupolar OP, as is often done for nematic ordering in liquid crystals. Consequently, our model draws parallels to the Landau-de Gennes theory of phase transitions in liquid crystals [29].
@@ II.2 Free-energy density
${\bm{\epsilon}}=\frac{\mathbb{1}}{3}\left[\alpha\left(T-T_{0}\right)-\frac{P}{K}+\left(\frac{e}{2\mu}-\frac{g}{K}\right)\operatorname{Tr}\bm{Q}^{2}\right]-\frac{d}{2\mu}\bm{Q}-\frac{e}{2\mu}\bm{Q}^{2}.$ (2)
@@
$\tilde{b}$ $=b-\frac{de}{2\mu},$
$\tilde{c}$ $=c-\frac{1}{2}\left(\frac{g^{2}}{K}+\frac{e^{2}}{12\mu}\right),$
$\tilde{T}_{Q}$ $=\left(T_{Q}+\frac{d^{2}}{4\mu a}+\frac{g\alpha}{a}T_{0}\right)\left(1+\frac{g\alpha}{a}\right)^{-1}.$